Martin is 5 years older than his sister. Five years ago, he was 2 times older than his sister. How old is Martin and his sister now? What is a system of equations for this question?
Answer by Timnewman(323)

Martin is 5 years older than his sister. Five years ago, he was 2 times older than his sister. How old is Martin and his sister now? What is a system of equations for this question?

Solution.
Let the sister's age be x yrs
then,
from your first sentence,
Martin's age =(x+5) yrs
From your second sentence,
Five years ago,
sister's age=x-5
then,
martin's age=(x+5)-5=x
but,martin was two times older,
therefore,
x=2(x-5)
solving the resulting equation,
x=2x-10
2x-x=10
x=10
sister's age is 10yrs,
martin's age=x+5
=10+5=15yrs